# DreamAudio: Customized Text-to-Audio Generation with Diffusion Models

### v1

@@ III-B1 Text Embedding
[…]
@@
where $\textit{f}_{\text{T5}}(\cdot)$ is a text encoder, chosen to be the FLAN-T5 model [53] in this work.
For each caption of the reference audio concepts, $\boldsymbol{t}_{k},k=1,...,K$, we first use the FLAN-T5 model to compute its embedding as $\boldsymbol{e}_{k}=\textit{f}_{\text{T5}}(\boldsymbol{t}_{k})$. We then concatenate the embeddings of all the reference captions as the second input for the module, as follows,
@@
This enables the model to interact with all the text embeddings of the referenced concepts.
@@ III-B2 Audio Embedding
We follow the baseline models [2] by applying a pre-trained VAE model to encode the reference audio clip from the mel-spectrogram into intermediate representations in a latent space. Similar to the text embedding for reference captions, the latent representation for each referenced audio $\boldsymbol{a}_{k}$ is obtained from the VAE encoder $\textit{f}_{\text{VAE}}(\cdot)$, as follows:
@@
The embeddings $\boldsymbol{r}_{k},k=1,...,K,$ are then concatenated as the third input:
@@
The VAE decoder is trained to turn this latent vector $\boldsymbol{r}_{k}$ to its original form $\boldsymbol{a}_{k}$, which is applied in the final stage in Section III-D to reconstruct the output waveforms.
Figure 3: The details of the MRC UNet network, where the model takes the reference feature $\boldsymbol{R}$ and $\boldsymbol{E}$, prompt feature $\boldsymbol{C}$ and the current noise data $\boldsymbol{z}_{\lambda}$ as inputs to generate the dynamics for denoised data $\boldsymbol{z}_{1}$ on step $\lambda$. The output $\mu(\cdot)$ can then be used for both training and inference.
@@ III-C Audio Feature Generation
III-C1 Module Strategy (Rectified Flow Matching)
Traditional diffusion-based generative models, such as DDPMs [11], generate samples through iterative denoising from a Gaussian prior, requiring hundreds of steps to gradually turn the noise distribution into the target data. However, such approaches are highly dependent on noise schedules and can be slow in sampling. Rectified Flow Matching (RFM) [74, 75], on the other hand, addresses the issue of slow sampling and increases stability by learning a continuous flow between noise and data distributions.
Unlike DDPMs that operate on discrete time steps $t\in\{0,1,...,N\}$, RFM models introduce a continuous flow variable $\lambda\in[0,1]$ that smoothly interpolates between Gaussian noise $\boldsymbol{z}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and the target data representation $\boldsymbol{z}_{1}$. By replacing discrete time steps with a continuous flow parameter $\lambda$, RFM enables smooth noise-to-data interpolation and a stable, schedule-free training objective based on a constant velocity field. The noisy data $\boldsymbol{z}_{\lambda}$ at the flow parameter of $\lambda$ is defined by a linear equation:
@@ III-C1 Module Strategy (Rectified Flow Matching)
To train the model, a neural network $\mu(\cdot)$, which will be introduced as MRC in the next section, is optimized to predict the velocity field $\boldsymbol{v}=\boldsymbol{z}_{1}-\beta\boldsymbol{z}_{0}$. Unlike denoising targets in general diffusion models, this velocity field is independent of $\lambda$, which provides a consistent supervision signal in all flow positions sampled and improves training stability. The network is trained to minimize the following objective:
@@
During inference, an ODE solver integrates the dynamics defined by the trained network $\mu(\cdot)$, starting from Gaussian noise $\boldsymbol{z}_{0}$ and producing $\boldsymbol{z}_{1}$. Compared to the hundreds or even thousands of iterations required in diffusion-based sampling [10], this process typically requires fewer than 50 integration steps to synthesize high-quality audio.
III-C2 Module Structure (Multi-Reference Customization)
To generate the customized audio feature with referenced concepts in a tuning-free manner, a multi-reference customization structure is designed for the generator module. As illustrated in Figure 3, the module integrates both text and audio feature from input prompt and reference concepts and predicts the dynamics to generate the denoised output $\boldsymbol{z}_{1}$. Our module is inspired by ControlNet [30], where a similar structure is designed to connect the reference feature and the generated outputs. In MRC, a UNet structure with new encoder blocks during down-sampling to handle referenced concepts, and external connections between new encoder blocks and decoder blocks to integrate features during up-sampling. Different from ControlNet, two groups of encoder blocks are designed with same structure and two groups of the reference feature are connected separately with decoder blocks and cross-attention blocks respectively.
@@ III-C2 Module Structure (Multi-Reference Customization)
Specifically, the proposed down-sampling process consists of two groups of encoder paths: feature encoding path for the generated feature vector and referenced feature extraction path for the reference feature vector. In the down-sampling stages, the encoder layers of the feature encoding take the noisy latent feature $\boldsymbol{z}_{\lambda}$ and the text embedding $\boldsymbol{C}$. The encoder layers of referenced feature extraction, on the other hand, take the latent representation of the referenced audio concept $\boldsymbol{R}$ and the corresponding textual embedding $\boldsymbol{E}$. This approach allows DreamAudio to identify and extract the audio feature for the customized content based on the reference audio and their corresponding captions. It is noted that two groups of encoder blocks share the same structure, e.g., convolutional layers with a cross-attention module for text embedding, but are applied with independent weights for different downsampling purposes, respectively.
@@
For the up-sampling stages, we feed all the features and conditions through a group of decoder blocks, designed as the feature decoding path. Specifically, the feature of the generated output and the extracted audio feature are integrated during the up-sampling stages. The reference audio features from each scaling stage are then considered as external conditions forwarded to the upsample layer within the same level. Different from single text embedding conditions in the feature encoding path, both the target prompt $\boldsymbol{E}$ and reference captions $\boldsymbol{R}$ are given as conditions through two cross-attention blocks, respectively.
@@ III-D Audio Feature Reconstruction
As shown in the structure of Figure 2, DreamAudio leverages both the VAE decoder and a GAN-based vocoder for reconstructing the audio feature from latent space into the target waveform. Followed by the RFM model, the generated audio representation is first decoded into the mel-spectrogram by the VAE decoder. In the next stage, a vocoder is applied to convert the audio feature into the waveform as the final output, and we train the vocoder using the SOTA structure BigVGAN [50] in the proposed system.
@@ III-E Customized Data Processing
In the early training stages, we observed that providing excessively detailed audio features caused the model to become overly reliant on these given representations, thereby limiting its ability to independently learn and generate new audio characteristics. To mitigate this, augmentation strategies are implemented to enhance the generalisation ability beyond provided references and improve robustness. Firstly, we randomly mask the contents of the customised audio event for the target audio. Secondly, we randomly remove the customised content during training. We train the DreamAudio model with a $10\%$ content masking and a $40\%$ content dropping rate. Experiments of the effectiveness of these masking and dropping metrics are introduced in Section VI-D3.
@@ IV Datasets and Evaluation Benchmark
Previous text-to-audio generation models [2, 52, 15] mainly work on public audio-language datasets [12, 13, 14], which contains audio clips and their corresponding captions. In customized generation tasks, additional reference concepts, such as audio clips and their corresponding captions are also required. To this end, we construct the customized dataset for training and evaluation of the proposed model based on four commonly used audio-language datasets. The construction of these datasets enables DreamAudio to extract the customized feature from specific user contents. In addition, we manually collect a small-scale dataset with real-world customized scenarios as the first benchmark for CTTA. In the following sections, we first introduce the datasets collected as the base dataset to form the customized datasets. Then, we discuss the strategies for developing the training and testing data, followed by the introduction of the benchmark dataset. All evaluation benchmarks used for the system are presented in the end.
@@
TABLE I: The setup of four datasets used for training and testing of the proposed DreamAudio.The reference-type indicates the strategy for generating the reference concept and the reference-num presents the number of audio events for customization through the reference concept.
@@ IV-A1 AudioCaps
AudioCaps (AC) [31] is one of the largest audio datasets with hand-crafted captions. As a subset of the AudioSet [12], AudioCaps contains $52,905$ $10$-second audio clips, each sample is matched with a human-annotated caption. Taking from original videos on YouTube, AudioCaps contains a variety of classes of audio, including music, Foley sounds [55] and human speech. Each audio clip has a single caption in the training set and five captions in the test set, where we only apply the first caption of each waveform to the testing data. We follow the official dataset split and collect $49,502$ audio clips to construct the training datasets and $928$ clips for building the testing datasets.
IV-A2 WavsCaps
WavCaps (WC) [13] is a machine-labeled dataset of audio captions generated through Large Language Models (LLMs). WavCaps contains $403,050$ audio clips collected from various datasets such as AudioSet [12] and FreeSound [76], providing various durations of audio clips ranging from $0.1$s seconds to $68$ seconds. To align the length of the input and target audio clips, we only collect the samples with length of less than $10$ seconds, resulting in a group of $163,818$ audio samples. Then, we randomly select $95\%$ of the data for training and $5\%$ for testing.
@@ IV-A3 UrbannSound8K
UrbanSound8K (UB8K) [77] contains $8,732$ labeled audio samples and each clip provides less than $4$ seconds of sound. The dataset is divided into $10$ different classes, including urban noise, background sound sources and natural sound sources. We randomly selected $8,000$ samples for training and $732$ for testing.
@@ IV-A4 ESC-50
ESC-50 (ESC) [78] dataset has $2,000$ 55-second audio recordings which are evenly categorized into $50$ distinct classes including natural sounds and domestic sounds. For the split of training and testing data, we randomly choose $1,600$ samples in ESC-50 for training and $400$ samples for testing.
@@ IV-B Customized Datasets
Based on the four general datasets mentioned in the previous section, we developed three different pipelines to generate customized datasets for training and evaluation. An overview of each dataset is presented in Table I and the pipelines for generating such datasets are illustrated in Figure 4.
@@ IV-B1 Customized-Concatenation
We develop the dataset by concatenating audio events. In detail, we collect short audio clips of less than $5$ seconds from the database and then randomly form a group of audio clips which lengths sum up to $10$ seconds. The target audio is then generated by concatenating these short audio clips. Next, we construct the referenced audio by grouping the short audio clips into three sets, each concatenated into a single clip as the reference audio. For the reference captions, we first transfer the labels from US8K and ESC-50 into short captions by simply adding verbs and subjects to provide more understandable captions. Then, we generate the sentence by connecting the caption of each short audio clip with a connection phrase randomly selected from our connection list. For the Customized-Concatenation dataset, we collect $92,300$ audio-text pairs for training and $200$ pairs for testing.
@@ IV-B2 Customized-Overlay
We find that the simple concatenating strategy led to several limitations. First, not all audio events can be connected smoothly, which may leave blank space. Second, audio created solely by concatenation often sounds unnatural, failing to accurately reproduce user-provided reference content, thus not effectively reflecting realistic scenarios. To guide the model with more realistic cases, we develop a second version of the customized dataset, called Customized-Overlay dataset, by mimicing real-world scenarios. Specifically, for each audio sample, we take a $10$-second audio clip as the base waveform and randomly select two shorter clips (less than $5$ seconds) as the “front” and “back” audio clips. The target audio clip is then generated by adding the “front” and “back” audio events into the base waveform. We remix the waveform of extended clips and base clips under a random signal to noise ratio (SNR) between -15 and 15 dB. The captioning strategy is similar to the approach we apply in the Customized-Concatenation dataset. We generated $146,481$ pairs for training and $200$ pairs for testing in the Customized-Overlay dataset.

### v2

@@ III-B1 Text Embedding
For both the target text prompts and the captions of the reference audio clips, we use the pre-trained Flan-T5 [60] as the text encoder to extract the text feature. Compared to contrastive language pretraining models, such as CLIP [59] and CLAP [84], the Flan-T5 encoder captures both semantic meaning [22] and temporal structures [87] from textual prompts, showing excellent performance in extracting semantic information for text-to-audio generation tasks [22, 91]. Denoting the text prompt as $\boldsymbol{c}$, the text embedding $\boldsymbol{C}$ is obtained as:
@@
where $\textit{f}_{\text{T5}}(\cdot)$ is the Flan-T5 text encoder [60].
For each caption of the reference audio concepts, $\boldsymbol{t}_{k},k=1,...,K$, we first use the same Flan-T5 model to compute the embedding as $\boldsymbol{e}_{k}=\textit{f}_{\text{T5}}(\boldsymbol{t}_{k})$. We then concatenate the embeddings of the reference captions as the reference embedding input $\boldsymbol{E}$, as follows:
@@
This enables the model to interact with all the text embeddings of the reference concepts.
@@ III-B2 Audio Embedding
We first follow the baseline models [43] by applying a pre-trained VAE encoder $\textit{f}_{\text{VAE}}(\cdot)$ to encode the reference audio clip from the mel-spectrogram into intermediate representations within the latent space. Taking $\boldsymbol{a}_{\lambda}$ as the mel-spectrogram of the original waveform $\mathbf{a}$, the latent feature for audio is obtained as $\boldsymbol{z}_{\lambda}=\textit{f}_{\text{VAE}}(\boldsymbol{a}_{\lambda})$. Similarly, the latent representation for the mel-spectrogram of referenced audio $\boldsymbol{a}_{k}$ is obtained from the VAE encoder [43], as follows:
@@
The embeddings $\boldsymbol{r}_{k},k=1,...,K,$ are then concatenated into $\boldsymbol{R}$, shown as input in Figure 3:
@@
Figure 3: The details of the MRC module, which takes the reference feature $\boldsymbol{R}$ and $\boldsymbol{E}$, prompt feature $\boldsymbol{C}$ and the current noisy data $\boldsymbol{z}_{\lambda}$ as inputs to generate the dynamics for denoised data $\boldsymbol{z}_{1}$ on step $\lambda$. The output $\mu(\cdot)$ can then be used for both training and inference.
@@ III-C Audio Feature Generation
III-C1 Rectified Flow Matching
Traditional diffusion-based generative models, such as DDPMs [26], generate samples through iterative denoising from a Gaussian prior, requiring hundreds of steps to gradually turn noise into the target data. However, such approaches are highly dependent on noise schedules and can be slow in sampling. RFM [48], on the other hand, addresses the sampling issue and increases stability by learning a continuous flow between noise and data distributions, as demonstrated in audio related tasks [92, 82].
Unlike DDPMs that operate on discrete time steps $t\in\{0,1,...,N\}$, RFM models introduce a continuous flow variable $\lambda\in[0,1]$ that smoothly interpolates between Gaussian noise $\boldsymbol{z}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and the target data representation $\boldsymbol{z}_{1}$. By replacing discrete time steps with a continuous flow parameter $\lambda$, RFM enables smooth noise-to-data interpolation and a stable, schedule-free training objective based on a constant velocity field $\boldsymbol{v}$. The noisy data $\boldsymbol{z}_{\lambda}$ at the flow parameter of $\lambda$ is defined by a linear equation:
@@ III-C1 Rectified Flow Matching
The model aims to estimate this velocity field $\boldsymbol{v}$, which defines the direction and magnitude of the continuous transformation that transports samples from the noise distribution to the data distribution. To train the model, a neural network $\mu(\cdot)$, corresponding to the proposed MRC module, is optimized to predict the velocity field $\boldsymbol{v}=\boldsymbol{z}_{1}-\beta\boldsymbol{z}_{0}$. Unlike denoising targets in general diffusion models, this velocity field is independent of $\lambda$, which provides a consistent supervision signal in all flow positions sampled and improves training stability. The network is trained to minimize the following objective:
@@
During inference, an ordinary differential equation (ODE) of the form $\frac{d\boldsymbol{z}_{\lambda}}{d\lambda}=\mu(\boldsymbol{z}_{\lambda},\boldsymbol{R},\lambda,\boldsymbol{E},\boldsymbol{C})$ is solved using the numerical ODE solver [48], starting from Gaussian noise $\boldsymbol{z}_{0}$ and integrating from $\lambda=0$ to $\lambda=1$ to obtain the final sample $\boldsymbol{z}_{1}$. Compared to the hundreds or even thousands of iterations required in diffusion-based sampling [74], this process typically requires fewer than 50 iterations to synthesize high-quality audio.
III-C2 Multi-Reference Customization
To produce customized audio features that incorporate referenced concepts without requiring fine-tuning, we designed the MRC structure. As shown in Figure 3, the MRC module merges textual and acoustic inputs, drawn from the input prompt and referenced concepts, and predicts the necessary dynamics to yield the denoised output $\boldsymbol{z}_{1}$.
The design draws inspiration from ControlNet [93], which adds a conditional encoder to steer the decoder. However, as illustrated in Figure 3, MRC adopts a U-Net backbone with dedicated encoder blocks in the down-sampling stage to process multiple referenced concepts. These new encoder blocks are externally linked to corresponding decoder blocks during up-sampling, enabling feature integration at multiple scales. Unlike ControlNet, which typically processes spatially aligned visual conditions (e.g., edges or depth maps), MRC handles multiple audio-text references with an external encoder. These references are fed into the generation pipeline via two separate mechanisms: extracted audio latent features are fused directly with the input of each decoder block for feature-level guidance, while text features are directed to cross-attention layers to ensure semantic alignment.
Rather than merely copying information from the references, MRC learns to identify relevant audio-event characteristics from multiple sources and blend them into the target generation. Specifically, the down-sampling stage comprises two parallel encoder paths: one for the generated feature vector and the other for the referenced feature vector. Their outputs are concatenated before being passed to a unified up-sampling decoder, ultimately producing the final latent representation.
Feature Encoding Path. The encoder layers take the noisy latent representation $\boldsymbol{z}_{\lambda}$ and the text prompt embedding $\boldsymbol{C}$ as input. Formally, for each encoding block $\mathcal{E}^{(l)}(\cdot)$, the latent feature is updated along with the flow parameter $\lambda$ as
$\boldsymbol{z}_{\lambda}^{(l)}=\mathcal{E}^{(l)}\!\left(\boldsymbol{z}_{\lambda}^{(l-1)},\;\lambda,\;\mathrm{A}\!\left(\boldsymbol{z}_{\lambda}^{(l-1)},\boldsymbol{C}\right)\right)$ (7)
where $\boldsymbol{z}_{\lambda}^{(0)}=\boldsymbol{z}_{\lambda}$ denotes the input noisy latent feature, $\mathcal{E}^{(l)}(\cdot)$ represents the $l$-th encoding block, and $\mathrm{A}(\cdot)$ denotes the attention operation that injects text conditioning information $\boldsymbol{C}$ into the latent representation.
@@ III-C2 Multi-Reference Customization
Reference Feature Extraction Path. The encoder layers for reference feature extraction, on the other hand, take the latent representation of the reference audio concept $\boldsymbol{R}$ and the corresponding textual embedding $\boldsymbol{E}$. Applying the same equation (7), this approach allows DreamAudio to identify and extract the audio feature for the customized content based on the reference audio and their corresponding captions. It is noted that the two groups of encoder blocks share the same structure, e.g., convolutional layers with a cross-attention module for text embedding, but are applied with independent weights for different downsampling purposes, respectively.
@@
Feature Decoding Path. We feed all the features and conditions through a group of decoder blocks, designed as the feature decoding path. Specifically, during the up-sampling stages, the latent vector from the feature encoding path and the feature extraction path are concatenated with the audio feature from each scaling level in downsampling stages via skipped connections, before passed into each decoder block. Different from single text embedding conditions in the feature encoding path, both the target prompt $\boldsymbol{C}$ and reference captions $\boldsymbol{E}$ are given as conditions through two cross-attention blocks respectively. The updated latent state for the $l$-th decoding block $\mathcal{D}^{(l)}$ is formed as:
$\boldsymbol{z}_{con}^{(l)}=\mathcal{D}^{(l)}\!\left(\boldsymbol{z}_{con}^{(l-1)},\lambda,\mathrm{A}_{1}\!\left(\boldsymbol{z}_{con}^{(l-1)},\boldsymbol{C}\right),\mathrm{A}_{2}\!\left(\boldsymbol{z}_{con}^{(l-1)},\boldsymbol{E}\right)\right)$
where $\boldsymbol{z}_{con}^{(l-1)}=[\boldsymbol{z}_{\lambda}^{(l-1)};\boldsymbol{R}^{(l-1)}]$ is the concatenation of the latent state from both the feature encoding path and the reference feature extraction path. $\mathrm{A}_{1}(\cdot)$ and $\mathrm{A}_{2}(\cdot)$ are the two blocks that calculate the cross-attention for the prompt embedding and the reference embedding, respectively.
@@ III-D Audio Feature Reconstruction
As shown in Figure 2, DreamAudio leverages both the VAE decoder and a generative adversarial network (GAN)-based vocoder for reconstructing the audio feature from latent space into the target waveform. Followed by the RFM model, the generated audio representation is first decoded into the mel-spectrogram by the VAE decoder. In the next stage, a vocoder is applied to convert the audio feature into the waveform as the final output, and we train the vocoder using the SOTA structure BigVGAN [42] in the proposed system.
@@ III-E Customized Data Processing
In the early training stages, we observed that providing excessively detailed audio features caused the model to become overly reliant on these given representations, thereby limiting its ability to independently learn and generate new audio characteristics. To mitigate this, augmentation strategies are implemented to enhance the generalisation ability beyond the provided references. First, we randomly mask the contents of the customised audio event for the target audio. Second, we randomly remove the customised content during training. We train the DreamAudio model with a $10\%$ content masking and a $40\%$ content dropping rate. These augmentation strategies, i.e. masking and dropping, are applied to both reference audio and reference text to improve the model’s robustness. Experiments on the effectiveness of these masking and dropping metrics are given in Section VI-D4.
@@ IV Datasets and Evaluation Benchmark
Previous text-to-audio generation models [43, 22, 88] mainly work on public audio-language datasets [21, 53, 76] These datasets only provide audio clips and their corresponding captions. In customized generation tasks, additional reference concepts, including reference audio clips and corresponding captions are also required. To this end, we construct the customized training and evaluation datasets based on four commonly used audio-language datasets. The construction of these datasets enables DreamAudio to extract the customized feature from specific user contents. In addition, we manually collect a small-scale dataset with real-world customized scenarios as the first benchmark for CTTA. In the following sections, we first introduce the datasets collected as the base dataset to form the customized datasets. Then, we discuss the strategies for developing the training and testing data, followed by the introduction of the benchmark dataset. All evaluation benchmarks used for the system are presented in the end.
@@
TABLE I: The setup of four datasets used for training and testing of the proposed DreamAudio. The reference-type indicates the strategy for generating the reference concept and the reference-num represents the number of audio events for customization through the reference concept.
@@ IV-A1 AudioCaps
AudioCaps (AC) [32] is one of the largest audio datasets with hand-crafted captions. As a subset of the AudioSet [21], AudioCaps contains $52,905$ $10$-second audio clips, and each clip is matched with a human-annotated caption. Taking from original videos on YouTube, AudioCaps contains various classes of audio, including music, Foley sounds [89], and human speech. Each audio clip has a single caption in the training set and five captions in the test set, where we only apply the first caption of each waveform to the testing data. We follow the official dataset split, and build the training set of $49,502$ audio clips and the testing set of $928$ clips.
IV-A2 WavCaps
WavCaps (WC) [53] is a machine-labeled dataset with audio captions generated through Large Language Models (LLMs). WavCaps contains $403,050$ audio clips collected from various datasets such as AudioSet [21] and FreeSound [17], providing various durations of audio clips ranging from $0.1$s seconds to $68$ seconds. To align the length of the input and target audio clips, we only collect the clips shorter than $10$ seconds, resulting in a group of $163,818$ audio clips. Then, we randomly selected $5\%$ for testing, forming a set of similar size to the AudioCaps test set.
@@ IV-A3 UrbannSound8K
UrbanSound8K (UB8K) [69] contains $8,732$ labeled audio clips and each clip is shorter than $4$ seconds. The dataset is divided into $10$ different classes, including urban noise, background sound sources, and natural sound sources. We randomly selected $732$ clips for model testing, similar to the size of the AudioCaps test set, and the remaining 8000 clips for model training.
@@ IV-A4 ESC-50
ESC-50 (ESC) [56] has $2,000$ 55-second audio recordings natural sounds and domestic sounds. These samples are evenly categorized into $50$ distinct classes. To match a similar data scale to AudioCaps, we randomly chose $400$ audio clips for model testing and $1,600$ clips for model training.
@@ IV-B Customized Datasets
With the four general audio datasets mentioned above, we design three task-specific data generation pipelines to construct customized datasets for training and evaluation. Unlike standard audio augmentation techniques, these pipelines are specifically designed for the CTTA task to simulate event-level customization. An overview of each customized dataset is presented in Table I, and the corresponding data generation pipelines are illustrated in Figure 4.
@@ IV-B1 Customized-Concatenation
We develop the dataset by concatenating audio events. In detail, we collect audio clips shorter than $5$ seconds from the database and then randomly form a group of audio clips whose lengths sum up to $10$ seconds. The target audio is then generated by concatenating these short audio clips. Next, we construct the referenced audio by grouping the short audio clips into three sets, each set is concatenated into a single clip as the reference audio. For the reference captions, we first convert the labels from US8K and ESC-50 into short captions by simply adding verbs and subjects. Then, we generate sentences by connecting the caption of each short audio clip with a connection phrase randomly selected from our connection list. For the Customized-Concatenation dataset, we collect $92,300$ audio-text pairs for model training and $200$ pairs for model testing.
@@ IV-B2 Customized-Overlay
We found that the simple concatenating strategy led to several limitations. First, not all audio events can be connected smoothly, which may leave a large blank space in the synthesized waveforms. Second, audio created solely by concatenation often sounds unnatural, failing to accurately reproduce user-provided reference contents, and thereby not reflecting realistic scenarios. To guide the model with more realistic cases, we develop a second version of the customized dataset, called Customized-Overlay dataset. Specifically, for each audio sample, we take a $10$-second audio clip as the base waveform and randomly select two shorter clips (less than $5$ seconds) as the “preceding” and “following” audio clips. The target audio clip is then generated by adding the “preceding” and “following” audio events into the base waveform. Then, we remix the waveform of two extended clips and the base clips under a random signal to noise ratio (SNR) between -15 and 15 dB. The captioning strategy is similar to the approach we apply in the Customized-Concatenation dataset. We generated $146,481$ pairs for training and $200$ pairs for testing in the Customized-Overlay dataset.